OpenStudy (anonymous):

if the sum of a number and five is tripled, the result is one less than twice the number. find the number.

OpenStudy (anonymous):

x=-16

OpenStudy (anonymous):

let the no. be = x sum of a number and five is tripled = 3(x + 5) 3(x + 5) = 2x - 1 3x + 15 = 2x -1 3x - 2x = -1 - 15 x = -16